40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 16 19:48:52 2016, Lydia, Update, General, Alignment status 
    Reply  Sat Sep 17 02:00:23 2016, rana, Update, SUS, Alignment status 
       Reply  Sat Sep 17 16:51:01 2016, Lydia, Update, SUS, Alignment status Screenshot_from_2016-09-17_16-45-00.png
          Reply  Sun Sep 18 16:18:05 2016, rana, Update, SUS, Alignment status 
             Reply  Mon Sep 19 11:11:43 2016, ericq, Update, SUS, Alignment status 
                Reply  Mon Sep 19 22:03:10 2016, ericq, Update, General, Further recovery progress 
                   Reply  Tue Sep 20 10:45:06 2016, rana, Update, General, Further recovery progress 
                      Reply  Wed Sep 21 01:08:02 2016, ericq, Update, General, Further recovery progress 
                         Reply  Wed Sep 21 09:04:57 2016, Steve, Update, General, 8 hours recovery progress 8hrs.png
                            Reply  Thu Sep 22 15:44:21 2016, Steve, Update, General, ITMX magnets are stucked again StuckAgain.png
                               Reply  Thu Sep 22 20:01:47 2016, Lydia, Update, General, ITMX freed, all optics kicked 
                                  Reply  Thu Sep 22 22:52:08 2016, ericq, Update, General, Damping found to be on 
Message ID: 12503     Entry time: Sun Sep 18 16:18:05 2016     In reply to: 12502     Reply to this: 12504
Author: rana 
Type: Update 
Category: SUS 
Subject: Alignment status 

susaux is responsible for turning on/off the inputs to the coil driver, but not the actual damping loops. So rebooting susaux only does the same as turning the watchdogs on/off so it shouldn't be a big issue.

Both before and after would be good. We want to see how much bias and how much voltage from the front ends were applied. l1susaux could have put in a huge bias, but NOT a huge force from the damping loops. But I've never seen it put in a huge bias and there's no way to prevent this anyway without disconnecting cables.

I think its much more likely that its a little stuck due to static charge on the rubber EQ stop tips and that we can shake it lose with the damping loops.

ELOG V3.1.3-